OpenStudy (anonymous):

Which expression is equivalent to sin2(π2−x)/sinx?

OpenStudy (anonymous):

A. –sin x B. sin x cot x C. cos x cot x D. sin x tan x

OpenStudy (campbell_st):

well what you need to know is \[\sin(\frac{\pi}{2} - x) = \cos(x)\] and then you need to know \[\sin^2(a) = 1 - \cos^2(a)\] you need these 2 properties to get to solution

OpenStudy (adryin):

\[\frac{ \sin2(\pi^{2}-x) }{ sinx }\]

OpenStudy (danjs):

i think you right assuming half pi there, so it goes to cos*cos*1/sin---